Mrs mitchell put a total of 10000 dls in two accounts. One account sarna 6% simple annual interest. The other earns 6.5% simple interest . After 1 year the two accounts earned 632.50 interest. Find how much money was invested in each account?
Answer by mananth(16946)

Part I 6.00% per annum ------------- Amount invested =x
Part II 6.50% per annum ------------ Amount invested = y
10000
Interest----- 632.50

Part I 6.00% per annum ---x
Part II 6.50% per annum ---y
Total investment
x + 1 y= 10000 -------------1
Interest on both investments
6.00% x + 6.50% y= 632.5
Multiply by 100
6 x + 6.5 y= 63250.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -60000.00
Add this to (2)
0 x 0.5 y= 3250
divide by 0.5
y = 6500
Part I 6.00% dls 3500
Part II 6.50% dls 6500

CHECK
3500 --------- 6.00% ------- 210.00
6500 ------------- 6.50% ------- 422.50
Total -------------------- 632.50

m.ananth@hotmail.ca